OpenStudy (anonymous):

Sonji paid $25.00 for two scarves, which were different prices. When she got home she could not find the receipt. She remembered that one scarf cost $3 more than the other. What was the price of the more expensive scarf? a. $11.00 b. $12.00 c. $13.00 d. $14.00

OpenStudy (muscrat123):

can i help?

OpenStudy (anonymous):

please!

OpenStudy (muscrat123):

ok. so what dont u understand?

OpenStudy (muscrat123):

 which equation is true? to find the two different prices. it gives u the more expensive price options, so u would subtract 3 from each. then take the new number and add it to the original number. which ever equation equals 25 is ur answer. (the equation that = 25's original number is the answer) tell me what u think/know the answer is .

OpenStudy (anonymous):

So it is 14?

OpenStudy (muscrat123):

correct! do u understand why?

OpenStudy (anonymous):

Let's represent scarf as "s". So \[\large s+(s+3)=25\]\[\large (s+s)=25-3\]\[\large 2s=22\]\[\large s=11\] So the scarf cost 11 dollars. Substitute that into the original equation to get \[\large s+(s+3)=25 \\\large 11+(11+3)=25\]

OpenStudy (anonymous):

I do now!!

OpenStudy (anonymous):

thanks

OpenStudy (muscrat123):

yw

OpenStudy (anonymous):

Woah woah woah. We're supposed to help solve the problem, not find the answer via the answer choices @muscrat123